### v1

# Dimension-Free Estimates for Noncommutative Discrete Maximal Spherical Means

Abstract.
In this paper, we establish dimension-free $L_{p}$-estimates for operator-valued maximal spherical means on cyclic groups ${\mathbb{Z}}_{m+1}^{d}$ for all $p>1$ and $m\geq 1$. The key ingredient is a noncommutative extension of the spectral technique developed by Nevo and Stein. As an application, we obtain a noncommutative spherical maximal inequality for automorphism actions of von Neumann algebras.
Key words and phrases:
Dimension-free bound, Noncommutative maximal inequalities, Noise operators, Noncommutative $L_{p}$-spaces
2020 Mathematics Subject Classification:
46L51, 42B20
1. Introduction
Let $B$ be a symmetric convex body in $d$-dimensional Euclidean space ${\mathbb{R}}^{d}$. For a locally integrable function $f$, the Hardy-Littlewood maximal function is defined by
$M_{B}f(v)=\sup_{t>0}\frac{1}{|B|}\int_{B}|f(v+tu)|du,$
where $|B|$ denotes the volume of $B$. For $1<p<\infty$, let $C_{p}(B)$ be the optimal constant such that
$\|M_{B}f\|_{p}\leq C_{p}(B)\|f\|_{p},\ \forall f\in L_{p}({\mathbb{R}}^{d});$
while $p=1$, $C_{1}(B)$ is taken to satisfy the weak type $(1,1)$ inequality,
$\|M_{B}f\|_{1,\infty}\leq C_{1}(B)\|f\|_{1},\ \forall f\in L_{1}({\mathbb{R}}^{d}).$
By applying the theory of spherical maximal functions, Stein [25] proved that for the $d$-dimensional Euclidean ball $B=B^{d}_{2}$ and for all $p>1$, the constant $C_{p}(B^{d}_{2})$ is bounded and independent of the dimension $d$. For $p=1$, Stein and Strömberg [26] later established the bound $C_{1}(B)\lesssim d\log d$. Subsequent work by Bourgain [2] and Carbery [4] extended these results, proving dimension-free bounds for $C_{p}(B)$ for general symmetric convex $B$ for $p>\frac{3}{2}$. For specific classes of symmetric convex bodies, such as the $d$-dimensional $\ell_{q}$-ball $B_{q}^{d}$, Müller [21] proved the dimension-free bounds for $C_{p}(B_{q}^{d})$ for all $p>1$ and $1\leq q<\infty$. In the case of the cube $B_{\infty}^{d}=[-1,1]^{d}$, Bourgain [3] showed that the dimension-free bounds of $C_{p}(B_{\infty}^{d})$ hold for all $1<p<\infty$. Furthermore, Aldaz [1] proved that $C_{1}(B_{\infty}^{d})$ grows to infinity as $d\to\infty$.
In recent years, dimension-free estimates have been extensively studied for discrete Hardy-Littlewood maximal operators. For $m\geq 1$, let ${\mathbb{Z}}^{d}_{m+1}$ denote the $d$-dimensional cyclic group equipped with the $\ell_{0}$ metric (Hamming distance),
$|u|=|\{1\leq i\leq d:u(i)\neq 0\}|$
for $u=(u(1),\cdot\cdot\cdot,u(d))\in{\mathbb{Z}}^{d}_{m+1}$. For each $1\leq k\leq d$, let
$\sigma_{k}:=\frac{1}{|\{|u|=k\}|}\chi_{\{|u|=k\}}=\frac{1}{\binom{d}{k}m^{k}}\chi_{\{|u|=k\}}$
denote the $L_{1}$-normalized indicator function of the $k$-sphere. The associated spherical maximal operator $M$ is given by
$M(f)(s)=\sup_{1\leq k\leq d}|\sigma_{k}\ast f(s)|,s\in{\mathbb{Z}}^{d}_{m+1}$
for $f:{\mathbb{Z}}^{d}_{m+1}\rightarrow{\mathbb{C}}$, where the convolution is defined as
$\sigma_{k}\ast f(s)=\sum_{u\in{\mathbb{Z}}^{d}_{m+1}}\sigma_{k}(u)f(s-u).$
The dimension-free boundedness of the maximal operator $M$ was first investigated by Harrow, Kolla and Schulman [8], who established its dimension-free $L_{2}$-boundedness on the hypercube ${\mathbb{Z}}^{d}_{2}$. Krause [18] later extended this result to all $p>1$. Subsequently, the dimension-free $L_{p}$-boundedness of $M$ was obtained for all cyclic groups ${\mathbb{Z}}^{d}_{m+1}$ with $m\geq 1$ in [14]. More precisely, Greenblatt, Kolla and Krause [14] proved that for any $p>1$ and any $m\geq 1$, there is a constant $C_{p,m}$ depends only on $p$ and $m$ such that
(1) $\|M(f)\|_{L_{p}({\mathbb{Z}}_{m+1}^{d})}\leq C_{p,m}\|f\|_{L_{p}({\mathbb{Z}}_{m+1}^{d})},\ \forall f\in L_{p}({\mathbb{Z}}_{m+1}^{d}).$
For recent advances in dimension-free estimates for $M$ on ${\mathbb{Z}}^{d}$, we refer the readers to [23] and the references therein.
In the area of noncommutative analysis, remarkable progress has been made in terms of maximal inequalities, for applications in noncommtative ergodic theory and noncommutative harmonic analysis. Notable examples include the noncommutative versions of maximal ergodic inequalities [17, 11, 6], Hardy-Littlewood maximal inequalities [20], Stein’s spherical maximal inequalities [9, 5, 19] etc. Despite the significant advances, the investigation of dimension-free bounds for noncommutative maximal inequalities remains largely unexplored, with only limited results in the literature.
The purpose of this paper is to establish the dimension-free $L_{p}$-bound (1) to the operator-valued setting, which then implies the corresponding noncommutative maximal inequalities. Let $\mathcal{M}$ be a von Neumann algebra equipped with a normal semifinite faithful trace $\tau$. The noncommutative $L_{p}$-spaces associated to the pair $({\mathcal{M}},\tau)$ is denoted by $L_{p}({\mathcal{M}})$. For each integer $k$, define the convolution operators
(2) $T_{k}f(s)=\sigma_{k}\ast f(s)=\frac{1}{\binom{d}{k}m^{k}}\sum_{u:|u|=k}f(s-u)\ ,$
for operator valued functions $f:{\mathbb{Z}}^{d}_{m+1}\rightarrow\mathcal{M}$. Let $\mathcal{N}=L_{\infty}({\mathbb{Z}}^{d}_{m+1},{\mathcal{M}})\cong L_{\infty}({\mathbb{Z}}^{d}_{m+1})\bar{\otimes}{\mathcal{M}}$ be the von Neumann algebra tensor product equipped with the tensor trace $\varphi=\mu\otimes\tau$, where $\mu$ is the Haar measure on ${\mathbb{Z}}^{d}_{m+1}$. Note that for $0<p<\infty$, $L_{p}(\mathcal{N})$ can be identified as the Bochner $L_{p}$-space $L_{p}({\mathbb{Z}}^{d}_{m+1};L_{p}({\mathcal{M}}))$. The following is our first main result.
Theorem 1.1.
Let $(T_{k})_{1\leq k\leq d}$ be the convolution operators defined as (2). Then for all $p>1$ and for all operator-valued functions $f\in L_{p}(\mathcal{N})\cong L_{p}({\mathbb{Z}}^{d}_{m+1};L_{p}({\mathcal{M}}))$
$\|(T_{k}f)_{1\leq k\leq d}\|_{L_{p}(\mathcal{N};\ell_{\infty})}\leq C_{p,m}\|f\|_{p},$
where constant $C_{p,m}$ depends only on $p$ and $m$ (independent of $d$).
Remark 1.2.
As noted in [14, Remark 1.6], the region $p>1$ is optimal because the optimal constant in the weak $(1,1)$ bound grows at least as fast as $\sqrt{d}$:
$\sup_{f\neq 0}\frac{\|Mf\|_{L_{1,\infty}({\mathbb{Z}}_{m+1}^{d})}}{\|f\|_{L_{1}({\mathbb{Z}}_{m+1}^{d})}}\geq c\sqrt{d}.$
An immediate difficulty in proving Theorem 1.1 is the fact that the classical maximal function $Mf=\sup_{1\leq k\leq d}|T_{k}f|$ does not make sense in general when $(T_{k}f)$ is a sequence of operators (see [17]). Indeed, the right formulation of noncommutative $L_{p}$-maximal inequalities was achieved via the introduction of vector-valued noncommutative $L_{p}$-spaces $L_{p}({\mathcal{M}};\ell_{\infty})$, which play the roles of the $L_{p}$-norm of maximal function $Mf$ in the noncommutative setting [16]. We refer to the next section for precise definitions.
As an application, we establish the noncommutative spherical maximal inequality for automorphism actions of ${\mathbb{Z}}_{m+1}^{d}$ on von Neumann algebras. Recall that for a von Neumann algebra $\mathcal{M}$}, we say that $\alpha:{\mathbb{Z}}_{m+1}^{d}\rightarrow\text{Aut}(\mathcal{M})$ is an action if for each $u$, $\alpha_{u}:\mathcal{M}\to\mathcal{M}$ is a *-preserving automorphism; and for all $s,u\in{\mathbb{Z}}_{m+1}^{d}$, $\alpha_{s}\circ\alpha_{u}=\alpha_{s+u}$. If in additional, $\tau\circ\alpha_{u}=\tau$ for all $u\in{\mathbb{Z}}_{m+1}^{d}$, we say $\alpha$ is a $\tau$-preserving action or an action of $\tau$-preserving automorphisms, denoted as $\alpha\curvearrowright(\mathcal{M},\tau)$. For each $k$, define
(3) $M_{k}x=\frac{1}{\binom{d}{k}m^{k}}\sum_{u:|u|=k}\alpha_{u}x,\ x\in{\mathcal{M}}.$
Theorem 1.1, together with the noncommutative Calderón transference principle developed in [11], implies the following result. Readers may refer to Section 2 for the definition of $\ell^{c}_{\infty}$-valued $L_{p}(\mathcal{M})$ space $L_{p}({\mathcal{M}};\ell^{c}_{\infty})$.
Theorem 1.3.
Let $\alpha$ be an $\tau$-preserving action of ${\mathbb{Z}}_{m+1}^{d}$ on a semi-finite von Neumann algebra $({\mathcal{M}},\tau)$. Let $(M_{k})_{1\leq k\leq d}$ be the defined as (3). Then for $p>1$,
$\|(M_{k}x)_{1\leq k\leq d}\|_{L_{p}({\mathcal{M}};\ell_{\infty})}\leq C_{p,m}\|x\|_{L_{p}({\mathcal{M}})},\ \forall x\in L_{p}({\mathcal{M}});$
and for $p>2$,
$\|(M_{k}x)_{1\leq k\leq d}\|_{L_{p}({\mathcal{M}};\ell^{c}_{\infty})}\leq\sqrt{C_{p,m}}\|x\|_{L_{p}({\mathcal{M}})},$
where the constant $C_{p,m}$ depends only on $p$ and $m$.
Theorem 1.3 can be rephrased as a maximal inequality for multiple ergodic averages. Let $(L_{i})_{1\leq i\leq d}$ be a commuting family (i.e. $L_{i}L_{j}=L_{j}L_{i}$) of $\tau$-preserving automorphisms on $\mathcal{M}$} such that $L_{i}^{m+1}=id$. For each positive integer $k$, define
(4) $N_{k}x=\frac{1}{\binom{d}{k}m^{k}}\sum_{n:|n|=k}L^{n}x,\ x\in{\mathcal{M}}.$
Here $n=(n_{1},\cdot\cdot\cdot,n_{d})$ and $L^{n}=L_{1}^{n_{1}}L_{2}^{n_{2}}\cdots L_{d}^{n_{d}}$. Then for $p>1$,
$\|(N_{k}x)_{1\leq k\leq d}\|_{L_{p}({\mathcal{M}};\ell_{\infty})}\leq C_{p,m}\|x\|_{L_{p}({\mathcal{M}})},\ \forall x\in L_{p}({\mathcal{M}});$
and $p>2$,
$\|(N_{k}x)_{1\leq k\leq d}\|_{L_{p}({\mathcal{M}};\ell^{c}_{\infty})}\leq\sqrt{C_{p,m}}\|x\|_{L_{p}({\mathcal{M}})},\ \forall x\in L_{p}({\mathcal{M}}).$
We conclude this introduction by outlining the structure of the paper. Section 2 reviews some basic background on noncommutative $L_{p}$-spaces. The proof of Theorem 1.1 are presented in Sections 3 and 4. We prove Theorem 1.3 in Section 5.
2. Preliminaries
Throughout this paper, we denote by $C$ a generic positive constant whose value may change between occurrences. The notation $X\lesssim Y$ indicates $X\leq CY$ for some constant $C>0$ independent of essential parameters, while $X\thickapprox Y$ signifies that both $X\lesssim Y$ and $Y\lesssim X$ hold simultaneously.
2.1. Noncommutative $L_{p}$-spaces
Let $\mathcal{M}$} be a von Neumann algebra equipped with a normal semi-finite faithful trace $\tau:{\mathcal{M}}_{+}\rightarrow[0,\infty]$. Let $\mathcal{S_{{\mathcal{M}}+}}$ be the set of all positive elements $x\in{\mathcal{M}}$ such that $\tau(\mathrm{supp}(x))<\infty$, where $\mathrm{supp}(x)$ denotes the support projection of $x$ and $\mathcal{S}_{{\mathcal{M}}}$ be the linear span of $\mathcal{S_{{\mathcal{M}}+}}$. Then $\mathcal{S}_{{\mathcal{M}}}$ forms a $w^{*}$-dense $\ast$-subalgebra of $\mathcal{M}$}. For $1\leq p<\infty$, the noncommutative $L_{p}$-space $L_{p}({\mathcal{M}},\tau)$ (in short, $L_{p}(\mathcal{M})$) is the completion of $\mathcal{S}_{{\mathcal{M}}}$ under the norm,
$\|x\|_{p}=[\tau(|x|^{p})]^{1/p}\ \ ,x\in\mathcal{S}_{{\mathcal{M}}},$
[…]
2.2. Vector-valued noncommutative $L_{p}$-spaces
We refer to [16, 17] for the references on vector-valued noncommutative $L_{p}$-spaces. Let $I$ be an index set. Given $1\leq p\leq\infty$, the space $L_{p}(\mathcal{M};\ell_{\infty}(I))$ consists of all families $(x_{n})_{n\in I}$ in $L_{p}(\mathcal{M})$ that admit a factorization $x_{n}=ay_{n}b$ with $a,b\in L_{2p}(\mathcal{M})$ and $(y_{n})_{n\in I}\subset\mathcal{M}$, equipped with the norm
$\|(x_{n})_{n\in I}\|_{L_{p}(\mathcal{M};\ell_{\infty}(I))}=\inf\left\{\big\|a\big\|_{{2p}}\sup_{n\in I}\big\|y_{n}\big\|_{\infty}\big\|b\big\|_{{2p}}\right\},$
where the infimum is taken over all possible factorizations $x_{n}=ay_{n}b,n\in I$. Following standard convention, this norm is often denoted concisely as $\|{\sup^{+}_{n\in I}}x_{n}\|_{p}$.
Remark 2.1.
[…]
$\|x\|_{L_{p}(\mathcal{M};\ell_{\infty}(I))}=\inf\Big\{\|a\|_{p}:a\in L^{+}_{p}({\mathcal{M}})\ \mbox{such that}\ x_{n}\leq a,\ \forall n\in I\Big\}.$
In the sequel, we omit the index set $I$ when it will not cause confusions.
As established in [16], for every $p>1$, the predual space of $L_{p}(\mathcal{M};\ell_{\infty})$ is $L_{p^{\prime}}(\mathcal{M};\ell_{1})$, where $p^{\prime}=\frac{p}{p-1}$ is the conjugate index of $p$. Given $1\leq p\leq\infty$, the space $L_{p}(\mathcal{M};\ell_{1})$ consists of all sequences $x=(x_{n})_{n}$ in $L_{p}(\mathcal{M})$ admitting a factorization of the form
$x_{n}=\sum_{k}u_{kn}^{\ast}v_{kn},\quad\forall n$
for two families $(u_{kn})_{k,n}$ and $(v_{kn})_{k,n}$ in $L_{2p}(\mathcal{M})$ such that
$\sum_{k,n}u^{\ast}_{kn}u_{kn}\in L_{p}(\mathcal{M})\;\;\text{and}\;\;\sum_{k,n}v^{\ast}_{kn}v_{kn}\in L_{p}(\mathcal{M}).$
The norm of $x$ in $L_{p}(\mathcal{M};\ell_{1})$ is defined by
$\|x\|_{L_{p}(\mathcal{M};\ell_{1})}=\inf\|\sum_{k,n}u^{\ast}_{kn}u_{kn}\|_{p}^{1/2}\|\sum_{k,n}v^{\ast}_{kn}v_{kn}\|_{p}^{1/2},$
where the infimum runs over all possible decompositions $x_{n}=\sum_{k}u_{kn}^{\ast}v_{kn}$. The duality between $L_{p}(\mathcal{M};\ell_{\infty})$ and $L_{p^{\prime}}(\mathcal{M};\ell_{1})$ is given by
$\langle x,\;y\rangle=\sum_{n}\tau(x_{n}y_{n}),\quad x=(x_{n})_{n}\in L_{p}(\mathcal{M};\ell_{\infty}),\;y=(y_{n})_{n}\in L_{p^{\prime}}(\mathcal{M};\ell_{1}).$
Remark 2.2.
A positive sequence $x=(x_{n})_{n}$ belongs to $L_{p}(\mathcal{M};\ell_{1})$ iff $\sum_{n}x_{n}\in L_{p}({\mathcal{M}})$; and in this case,
$\|x\|_{L_{p}(\mathcal{M};\ell_{1})}=\big\|\sum_{n}x_{n}\big\|_{p}.$
We collect some important properties of these two vector-valued noncommutative $L_{p}$-spaces. See [17] for the detailed proofs.
Proposition 2.3.
The following statements hold.
  • (i)
    Each element in the unit ball of $L_{p}({\mathcal{M}};\ell_{\infty})$ (resp. $L_{p}({\mathcal{M}};\ell_{1})$) can be written as a sum of sixteen (resp. eight) positive elements in the same ball.
  • (ii)
    Let $x=(x_{n})$ be a positive sequence in $L_{p}({\mathcal{M}};\ell_{\infty})$. Then for any $1\leq p\leq\infty$,
    $\|{\sup_{n}}^{+}x_{n}\|_{p}=\sup\Big\{\sum_{n}\tau(x_{n}y_{n}):y_{n}\in L_{p^{\prime}}({\mathcal{M}})\ \mbox{and}\ \big\|\sum_{n}y_{n}\big\|_{p^{\prime}}\leq 1\Big\}.$
    In particular, for two positive sequences $(x_{n})\leq(y_{n})$, $\|{\sup_{n}}^{+}x_{n}\|_{p}\leq\|{\sup_{n}}^{+}y_{n}\|_{p}.$
  • (iii)
    Let $x=(x_{n})$ be a sequence in $L_{p}({\mathcal{M}};\ell_{\infty})$ and $(z_{n,k})_{n,k}\in{\mathbb{C}}$. Then for any $1\leq p\leq\infty$,
    $\big\|{\sup_{n}}^{+}\sum_{k}z_{n,k}x_{k}\big\|_{p}\leq\sup_{n}(\sum_{k}|z_{n,k}|)\|{\sup_{n}}^{+}x_{n}\|_{p}.$
  • (iv)
    Let $1\leq p_{0}<p_{1}\leq\infty$ and $0<\theta<1$. We have the following complex interpolation relation,
    $L_{p}({\mathcal{M}};\ell_{1})=\big(L_{p_{0}}({\mathcal{M}};\ell_{1}),L_{p_{1}}({\mathcal{M}};\ell_{1})\big)_{\theta}\ ,\ \text{isometrically}$
    where $\frac{1}{p}=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}}$.
We then turn to the definition of $L_{p}({\mathcal{M}};\ell^{c}_{\infty})$. Let $2\leq p\leq\infty$, the space $L_{p}({\mathcal{M}};\ell^{c}_{\infty})$ is defined as the family of all $(x_{n})_{n}$ in $L_{p}({\mathcal{M}})$ for which there are an $a\in L_{p}({\mathcal{M}})$ and $(z_{n})_{n}\subseteq L_{\infty}({\mathcal{M}})$ such that
$x_{n}=z_{n}a\ \ \mbox{and}\ \ \sup_{n}\|z_{n}\|_{\infty}<\infty.$
The norm $\|(x_{n})_{n}\|_{L_{p}({\mathcal{M}};\ell^{c}_{\infty})}$ is then defined as to be the infimum $\inf\{\sup_{n}\|z_{n}\|_{\infty}\|a\|_{p}\}$ over all factorizations of $(x_{n})_{n}=(z_{n})a$ as above. Note that $(x_{n})_{n}\in L_{p}({\mathcal{M}};\ell^{c}_{\infty})$ iff $(x^{*}_{n}x_{n})_{n}\in L_{p/2}({\mathcal{M}};\ell_{\infty})$.
Finally, we introduce the $\ell_{\infty}$-valued noncommutative weak $L_{p}$-spaces $\Lambda_{p,\infty}({\mathcal{M}};\ell_{\infty})$ (see e.g. [10]). Given a sequence $(x_{n})_{n}$ in $L_{p}(\mathcal{M})$ with $1\leq p<\infty$, we define
$\|(x_{n})_{n}\|_{\Lambda_{p,\infty}({\mathcal{M}};\ell_{\infty})}=\sup_{\lambda>0}\ \lambda\inf\Big\{\big(\tau(e^{\perp})\big)^{\frac{1}{p}}:e\in\mathcal{P}({\mathcal{M}})\text{ s.t. }\|ex_{k}e\|_{\infty}\leq\lambda\ \mbox{for\ all}\ k\Big\},$
where $\mathcal{P}({\mathcal{M}})$ is the set of all projections in $\mathcal{M}$}. The space $\Lambda_{p,\infty}({\mathcal{M}};\ell_{\infty})$ consists of all sequences $x=(x_{n})_{n}$ for which this quasi-norm is finite.
2.3. Noise operators
Recall that the dual group of ${\mathbb{Z}}_{m+1}^{d}$ is itself ${\mathbb{Z}}_{m+1}^{d}$. In the following, we will denote points in the group by lowercase letters, and frequencies in the dual group by capital letters. For each $S,u\in{\mathbb{Z}}_{m+1}^{d}$, we define the $L_{2}$-normalized character
$\chi_{S}(u)=\frac{1}{(m+1)^{d/2}}\xi_{m}^{S\cdot u}=\frac{1}{(m+1)^{d/2}}\prod_{i=1}^{d}\xi_{m}^{S(i)u(i)},$
where $\xi_{m}=e^{2\pi i/(m+1)}$ is a primitive $(m+1)$-th root of unity and $S\cdot u=\sum_{i=1}^{d}S(i)u(i)$. A operator-valued function $f:{\mathbb{Z}}_{m+1}^{d}\rightarrow{\mathcal{M}}$ admits a Fourier expansion of the form:
$f(u)=\sum_{S\in{\mathbb{Z}}_{m+1}^{d}}\widehat{f}(S)\chi_{S}(u),$
where each Fourier coefficient is given by
$\widehat{f}(S)=\sum_{u\in{\mathbb{Z}}_{m+1}^{d}}f(u)\overline{\chi_{S}(u)}.$
For each $t>0$, the noise operator $N_{t}$ is defined by via its Fourier expansion
(5) $N_{t}f(u)=\sum_{S\in{\mathbb{Z}}_{m+1}^{d}}e^{-t|S|}\widehat{f}(S)\chi_{S}(u),$
and thus the family of operators $(N_{t})_{t\geq 0}$ form a semigroup. For each $T>0$, define
(6) $H_{T}f=\frac{1}{T}\int_{0}^{T}N_{t}fdt.$
The boundedness theory for the family $(H_{T})_{T\geq 0}$ was first established by Lance [15] for the case $p=\infty$, and later by Yeadon [27] for $p=1$. However, the case $1<p<\infty$ remained open for thirty years until it was finally resolved by Junge and Xu [17] through the development of noncommutative Marcinkiewicz interpolation. Junge-Xu’s result has since been widely applied in noncommutative ergodic theory and noncommutative harmonic analysis. We state their results as the following theorem.
Theorem 2.4 ([15, 27, 17]).
Let $\mathcal{N}=L_{\infty}({\mathbb{Z}}^{d}_{m+1})\bar{\otimes}{\mathcal{M}}$. The following assertions hold with a positive constant $C_{p,m}$ depending only on $p$ and $m$:
  • (i)
    for $p=1$ and any operator-valued function $f\in L_{1}(\mathcal{N})$,
    $\|(H_{T}f)_{T>0}\|_{\Lambda_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{L_{1}(\mathcal{N})};$
  • (ii)
    for $1<p\leq\infty$ and any operator-valued function $f\in L_{p}(\mathcal{N})$,
    $\|(H_{T}f)_{T>0}\|_{L_{p}(\mathcal{N};\ell_{\infty})}\leq C_{p,m}\|f\|_{L_{p}(\mathcal{N})}\ \ .$
3. The smooth spherical maximal operators
In this section, we study the boundedness theory of two smooth spherical maximal operators. By triangle’s inequality in $L_{p}(\mathcal{N};\ell_{\infty})$,
$\|(T_{k}f)_{1\leq k\leq d}\|_{L_{p}(\mathcal{N};\ell_{\infty})}$ $\leq\|(T_{k}f)_{1\leq k\leq\frac{md}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})}+\|(T_{k}f)_{\frac{md}{m+1}\leq k\leq d}\|_{L_{p}(\mathcal{N};\ell_{\infty})}$
$=\|(T_{k}f)_{1\leq k\leq\frac{md}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})}+\|(T_{d-k}f)_{1\leq k\leq\frac{d}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})},$
where the convolution operators $T_{k},1\leq k\leq d$ are defined in (2). Therefore, to establish Theorem 1.1, it suffices to estimate the aforementioned local term $(T_{k}f)_{1\leq k\leq\frac{md}{m+1}}$ and distant terms $(T_{d-k}f)_{1\leq k\leq\frac{d}{m+1}}$. For this purpose, we introduce the corresponding regularized operators for both local and distant terms. For $K\geq 0$, define
$T_{K}^{M}:=\frac{1}{K+1}\sum_{k\leq K}T_{k}\ ,\ T_{K}^{M}f=\frac{1}{K+1}\sum_{k\leq K}\sigma_{k}\ast f,$
$T_{K}^{D}:=\frac{1}{K+1}\sum_{k\leq K}T_{d-k}\ ,\ T_{K}^{D}f=\frac{1}{K+1}\sum_{k\leq K}\sigma_{d-k}\ast f.$
Proposition 3.1.
Let $1\leq p\leq\infty$. There exists a constant $C_{p,m}$ depending only on $p$ and $m$ such that
  • (i)
    for $p=1$ and any $f\in L_{1}(\mathcal{N})=L_{1}(\mathbb{Z}_{m+1}^{d};L_{1}(\mathcal{M}))$,
    $\|(T_{K}^{M}f)_{0\leq K\leq\frac{md}{m+1}}\|_{\Lambda_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{L_{1}(\mathcal{N})}\ ,$
    $\|(T_{K}^{D}f)_{0\leq K\leq\frac{d}{m+1}}\|_{\Lambda_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{L_{1}(\mathcal{N})}\ ;$
  • (ii)
    for $1<p\leq\infty$ and any $f\in L_{p}(\mathcal{N})=L_{p}(\mathbb{Z}_{m+1}^{d};L_{p}(\mathcal{M}))$,
    $\|(T_{K}^{M}f)_{0\leq K\leq\frac{md}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})}\leq C_{p,m}\|f\|_{L_{p}(\mathcal{N})}\ ,$
    $\|(T_{K}^{D}f)_{0\leq K\leq\frac{d}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})}\leq C_{p,m}\|f\|_{L_{p}(\mathcal{N})}\ .$
The proof of Proposition 3.1 relys on Theorem 2.4. To see this, given any fixed parameter $\eta\in(0,1)$, we define a probability measure $\mu_{\eta}$ on ${\mathbb{Z}}_{m+1}$ by $\mu_{\eta}(w)=1-\eta$ if $w=0$ and $\mu_{\eta}(w)=\frac{\eta}{m}$ otherwise. This induces the product measure on ${\mathbb{Z}}_{m+1}^{d}$: for $u\in{\mathbb{Z}}_{m+1}^{d}$,
$\mu^{d}_{\eta}(u)=\Big(\frac{\eta}{m}\Big)^{|u|}(1-\eta)^{d-|u|}.$
It was proved from [14, Lemma 4.6] that
(7) $\hat{\mu}_{\eta}(S)=\Big(1-\frac{(m+1)\eta}{m}\Big)^{|S|}\ ,\ f\ast\mu^{d}_{\eta}(u)=\sum_{S\in{\mathbb{Z}}_{m+1}^{d}}\Big(1-\frac{(m+1)\eta}{m}\Big)^{|S|}\widehat{f}(S)\chi_{S}(u).$
where $\hat{\mu}^{d}_{\eta}$ denotes the Fourier transform of $\mu^{d}_{\eta}$. For $0<P\leq\frac{m}{m+1}$, denote
(8) $\widetilde{N}_{\eta}f=f\ast\mu^{d}_{\eta}\ ,\ J_{P}f:=\frac{1}{P}\int_{0}^{P}\widetilde{N}_{\eta}fd\eta.$
Note that $\widetilde{N}_{\eta}=N_{(1-\frac{(m+1)\eta}{m})}$ is a reparametrization of the noise operator.
Lemma 3.2.
There exists a constant $C_{p,m}$ depending only on $p$ and $m$ such that
  • (i)
    for $p=1$,
    $\|(J_{P}f)_{0<P\leq\frac{m}{m+1}}\|_{\Lambda_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{1}\ \ \forall f\in L_{1}(\mathcal{N});$
  • (ii)
    for $1<p\leq\infty$,
    $\|(J_{P}f)_{0<P\leq\frac{m}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})}\leq C_{p,m}\|f\|_{p}\ \ \forall f\in L_{p}(\mathcal{N}).$
Proof.
In the proof, we may assume that $f$ is positive. Indeed, by decomposing $f=f_{1}-f_{2}+i(f_{3}-f_{4})$ with positive $f_{j}$ and $\|f_{j}\|_{p}\leq\|f\|_{p}$ for $j\in\{1,2,3,4\}$ and all $p\geq 1$, we can use (quasi-)triangle’s inequality to restrict the proof to the case of positive operators.
(i) Following [14, Proposition 4.7], we use the short notation $\rho_{m}:=\frac{m}{m+1}$ and define the measure
$\nu_{P}=\Big\{\frac{\rho_{m}}{P}Te^{-T}\chi_{(0,-\ln[1-P/\rho_{m}])}\Big\}dT+\Big\{\Big(\frac{\rho_{m}}{P}-1\Big)\Big(-\ln\Big[1-\frac{P}{\rho_{m}}\Big]\Big)\Big\}\delta_{-\ln[1-P/\rho_{m}]},$
for $0<P\leq\rho_{m}$, where $\delta$ denotes the dirac measure. One can verify that $\nu_{P}$ has total mass 1. Moreover, it was shown in [14, Proposition 4.7] that
(9) $J_{P}f=\int_{0}^{\infty}\Big(\frac{1}{T}\int_{0}^{T}N_{t}fdt\Big)d\nu_{P}(T)=\int_{0}^{\infty}H_{T}fd\nu_{P}(T).$
We now appeal to the weak type $(1,1)$ boundedness of $(H_{T})_{T}$ stated in Theorem 2.4 (i) that for any $\lambda>0$, there exists a projection $e\in\mathcal{N}$ such that $\mu\otimes\tau(1-e)\leq C_{1,m}\frac{\|f\|_{1}}{\lambda}$ and
$\sup_{0<P\leq\frac{m}{m+1}}\|eJ_{P}fe\|_{\infty}\leq\int_{0}^{\infty}\sup_{T>0}\|eH_{T}fe\|_{\infty}d\nu_{P}(T)\leq\lambda.$
This proves (i).
(ii) The strong type $(p,p)$ estimates follows analogously. Indeed, by Theorem 2.4 (ii) and Remark 2.1, there exist $F\in L^{+}_{p}({\mathcal{N}})$ and a constant $C_{p,m}$ such that for all $T>0$,
$H_{T}f\leq F\ \ \mbox{and}\ \ \|F\|_{p}\leq C_{p,m}\|f\|_{p}.$
It then follows from (9) that for all $0<P\leq\rho_{m}$,
$J_{P}f\leq\int_{0}^{\infty}Fd\nu_{P}(T)=F.$
Therefore, by using Remark 2.1 once more, we obtain
$\|\underset{0<P\leq\frac{m}{m+1}}{\text{sup}^{+}}J_{P}f\|_{p}\leq\|F\|_{p}\leq C_{p,m}\|f\|_{p},$
which finishes the proof. ∎
Now we are ready to prove Proposition 3.1 .
Proof of Proposition 3.1..
Without loss of generality, we can assume that $f$ is positive. By the kernel estimate presented in [14, Proposition 4.8], for each $K\leq d$, one can choose $P(K)\in(0,\frac{m}{m+1}]$ and an absolute constant $C>0$ such that
$\frac{1}{K+1}\sum_{l\leq K}\sigma_{l}\leq C\frac{1}{P(K)}\int_{0}^{P(K)}\mu^{d}_{\eta}d\eta.$
As a consequence,
$T_{K}^{M}f=\frac{1}{K+1}\sum_{l\leq K}\sigma_{l}\ast f\leq C\frac{1}{P(K)}\int_{0}^{P(K)}\widetilde{N}_{\eta}fd\eta=CJ_{P(K)}f.$
The desired bound for $T_{K}^{M}f$ then follows from the corresponding estimates in Lemma 3.2.
For $T^{D}_{K}$, by similar kernel estimate in [14, Proposition 4.8], we have that for any $0\leq L\leq\frac{md}{m+1}$, there exists a constant $C_{m}$ such that
$\frac{1}{\lfloor\frac{L}{m}\rfloor+1}\sum_{l\leq L/m}\sigma_{d-l}\leq C_{m}\frac{1}{L+1}\sum_{l\leq L}\sigma_{l}\ast\sigma_{d},$
where $\lfloor a\rfloor$ denotes the integer part of $a$. Hence, for any $K\leq\frac{d}{m+1}$
(10) $T_{K}^{D}f\leq C_{m}\frac{1}{L+1}\sum_{l\leq L}\sigma_{l}\ast(\sigma_{d}\ast f).$
By the weak $(1,1)$-estimate of $T_{K}^{M}$, for any $\lambda>0$ there exists a projection $e\in\mathcal{N}$ such that
$\sup_{0\leq K\leq\frac{d}{m+1}}\|e(T_{K}^{D}f)e\|_{\infty}\leq C_{m}\sup_{L\leq\frac{md}{m+1}}\|e(T^{M}_{L}(\sigma_{d}\ast f))e\|_{\infty}\leq C_{m}\lambda$
and
$\mu\otimes\tau(1-e)\leq\frac{\|\sigma_{d}\ast f\|_{1}}{\lambda}\leq C_{m}\frac{\|f\|_{1}}{\lambda},$
where we used the Young inequality ($\sigma_{d}$ is a probability measure)
(11) $\|\sigma_{d}\ast f\|_{p}\leq\|f\|_{p},\,\ \,\forall\ 1\leq p\leq\infty.$
Then we combine (10), the strong type $(p,p)$ estimate for $T_{K}^{M}$ and (11) to obtain
$\big\|\underset{0\leq K\leq\frac{d}{m+1}}{\text{sup}^{+}}T_{K}^{D}f\big\|_{p}$ $\leq C_{m}\big\|\underset{0\leq L\leq\frac{md}{m+1}}{\text{sup}^{+}}T^{M}_{L}(\sigma_{d}\ast f)\big\|_{p}$
$\leq C_{p,m}\|\sigma_{d}\ast f\|_{p}\leq C_{p,m}\|f\|_{p}.$
This proves the strong type $(p,p)$ estimates for $T_{K}^{D}$. ∎
4. Proof of Theorem 1.1
This section is devoted to the proof of Theorem 1.1. Let us begin by recalling the decomposition from the previous section:
$\|(T_{k}f)_{1\leq k\leq d}\|_{L_{p}(\mathcal{N};\ell_{\infty})}\leq\|(T_{k}f)_{1\leq k\leq\frac{md}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})}+\|(T_{d-k}f)_{1\leq k\leq\frac{d}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})}.$
Then it suffices to establish the following two estimates:
(12) $\|(T_{k}f)_{1\leq k\leq\frac{md}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})}\leq C_{p,m}\|f\|_{p};$
(13) $\|(T_{d-k}f)_{1\leq k\leq\frac{d}{m+1}}\|_{L_{p}(\mathcal{N};\ell_{\infty})}\leq C_{p,m}\|f\|_{p}.$
The proofs of (12) and (13) are inspired by the complex Cesaro sums method developed by Nevo and Stein [22]. We introduce some necessary notations. Given a complex number $\alpha$ and a nonnegative integer $n$, set
$A^{\alpha}_{n}=\frac{(\alpha+1)(\alpha+2)\dots(\alpha+n)}{n!},\ \ A_{0}^{\alpha}:=1,\ A_{-1}^{\alpha}:=0.$
The following elementary properties on $A_{n}^{\alpha}$ are well-known, see [22, Lemma 3].
Lemma 4.1.
Let $\alpha=\beta+i\gamma$ be a complex number with $\beta,\gamma\in{\mathbb{R}}$. There exists $c_{\beta}>0$ such that
  • (i)
    for $\beta>-1$,
    $c_{\beta}^{-1}(n+1)^{\beta}\leq A_{n}^{\beta}\leq c_{\beta}(n+1)^{\beta};$
  • (ii)
    for $\beta>-1$,
    $|A_{n}^{\beta}|\leq|A_{n}^{\alpha}|\leq c_{\beta}e^{2\gamma^{2}}|A_{n}^{\beta}|;$
  • (iii)
    for $\beta\in{\mathbb{N}}$,
    $|(1+n)^{\beta}A_{n}^{-\beta+i\gamma}|\leq c_{\beta}e^{3\gamma^{2}}.$
For $\alpha\in{\mathbb{C}}$ and $n\in{\mathbb{N}}$, we define
$S_{n}^{\alpha}f=\sum_{k=0}^{n}A_{n-k}^{\alpha}T_{k}f\ ,\ M_{n}^{\alpha}f=\frac{1}{(n+1)^{\alpha+1}}S_{n}^{\alpha}f;$
$U_{n}^{\alpha}f=\sum_{k=0}^{n}A_{n-k}^{\alpha}T_{d-k}f\ ,\ N_{n}^{\alpha}f=\frac{1}{(n+1)^{\alpha+1}}U_{n}^{\alpha}f.$
We will establish the following more general estimate.
Theorem 4.2.
Let $(\mathcal{M},\tau)$ be a semifinite von Neumann algebra. For any $\alpha\in{\mathbb{C}}$ and $p>1$, there exists a constant $C_{m,p,\alpha}$ depends only on $m,p,\alpha$ such that $f\in L_{p}(\mathcal{N})=L_{p}(\mathbb{Z}_{m+1}^{d};L_{p}(\mathcal{M}))$
$\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}M_{n}^{\alpha}(f)\|_{p}\leq C_{m,p,\alpha}\|f\|_{p};$
$\|{\underset{0\leq n\leq\frac{d}{m+1}}{\mathrm{sup}^{+}}}N_{n}^{\alpha}(f)\|_{p}\leq C_{m,p,\alpha}\|f\|_{p}.$
Indeed, Theorem 4.2 directly yields estimates (12) and (13) through the identifications $M_{n}^{-1}f=T_{n}f$ and $N_{n}^{-1}f=T_{d-n}f$. Our proof strategy for Theorem 4.2 follows the framework established by Nevo and Stein in [22, Section 3], which we adapt to the noncommutative setting. The proof of Theorem 4.2 will be divided into the following four lemmas.
Lemma 4.3.
Let $\alpha=\beta+i\gamma$ with $\beta\geq 0$. Then for all $p>1$ and $f\in L_{p}(\mathcal{N})$,
$\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}M_{n}^{\alpha}(f)\|_{p}\leq C_{m,p,\beta}e^{2\gamma^{2}}\|f\|_{p};$
$\|{\underset{0\leq n\leq\frac{d}{m+1}}{\mathrm{sup}^{+}}}N_{n}^{\alpha}(f)\|_{p}\leq C_{m,p,\beta}e^{2\gamma^{2}}\|f\|_{p}.$
Proof.
Without loss of generality, we assume that $f$ is positive. Let $(g_{n})\subset L^{+}_{p^{\prime}}(\mathcal{N})$ with $\|\sum_{n}g_{n}\|_{L_{p^{\prime}}(\mathcal{N})}\leq 1$. By applying Lemma 4.1, we have
$|\tau(M_{n}^{\alpha}(f)g_{n})|$ $\leq(n+1)^{-\beta-1}\sum_{k=0}^{n}|A_{n-k}^{\alpha}|\tau(T_{k}fg_{n})$
$\leq c_{\beta}e^{2\gamma^{2}}(n+1)^{-\beta-1}\sum_{k=0}^{n}(n-k+1)^{\beta}\tau(T_{k}fg_{n})$
$\leq c_{\beta}e^{2\gamma^{2}}(n+1)^{-1}\sum_{k=0}^{n}\tau(T_{k}fg_{n}).$
Therefore, recalling that $T_{n}^{M}f=(n+1)^{-1}\sum_{k\leq n}T_{k}f$, we obtain
$\sum_{0\leq n\leq\frac{md}{m+1}}|\tau(M_{n}^{\alpha}(f)g_{n})|\leq c_{\beta}e^{2\gamma^{2}}\sum_{0\leq n\leq\frac{md}{m+1}}\tau(T_{n}^{M}fg_{n}).$
Taking the supremum over all $(g_{n})\subset L^{+}_{p^{\prime}}(\mathcal{N})$ with $\|\sum_{n}g_{n}\|_{L_{p^{\prime}}(\mathcal{N})}\leq 1$, using Proposition 2.3 (i)(ii) and Proposition 3.1 (ii), we deduce the first assertion. The second statement can be proved similarly by using the definition of $T_{n}^{D}f$ and Proposition 3.1 (ii). ∎
The following $L_{2}$-estimates are identical to the scalar-valued cases in [14, Proposition 5.7], so we omit the proof. For an operator $x$, recall that $|x|=(x^{*}x)^{\frac{1}{2}}$.
Lemma 4.4.
Let $t\in{\mathbb{N}}$. Then there exists constant $C_{t,m}>0$ depending only on $t$ and $m$ such that for all $f\in L^{+}_{2}(\mathcal{N})$,
$\|R^{1}_{t,m}f\|_{2}\leq C_{t,m}\|f\|_{2}\ \mbox{ and }\ \|R^{2}_{t,m}f\|_{2}\leq C_{t,m}\|f\|_{2},$
where $R^{1}_{t,m}f$ and $R^{2}_{t,m}f$ are respectively defined by
$R^{1}_{t,m}f=\Big(\sum_{0\leq k\leq\frac{md}{m+1}}(k+1)^{2t-1}|S_{k}^{-t-1}f|^{2}\Big)^{\frac{1}{2}};$
$R^{2}_{t,m}f=\Big(\sum_{0\leq k\leq\frac{d}{m+1}}(k+1)^{2t-1}|U_{k}^{-t-1}f|^{2}\Big)^{\frac{1}{2}}.$
With Lemma 4.4 in hand, the next step is to prove the maximal inequalities in $L_{2}(\mathcal{N})$ for $M_{n}^{-t}$ and $N_{n}^{-t}$ with $t\in{\mathbb{N}}$. For positive integers $t$ and $k$, set
$B_{k}^{t}=k(k-1)\cdots(k-t+1)\ \mbox{for}\;t\leq k\quad\mbox{and}\quad B_{k}^{t}=0\ \mbox{for}\;t>k.$
Lemma 4.5.
Let $t\in{\mathbb{N}}$. Then there exists $C_{t,m}$ such that for all $f\in L_{2}(\mathcal{N})$
$\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}M_{n}^{-t}(f)\|_{2}\leq C_{t,m}\|f\|_{2};$
$\|{\underset{0\leq n\leq\frac{d}{m+1}}{\mathrm{sup}^{+}}}N_{n}^{-t}(f)\|_{2}\leq C_{t,m}\|f\|_{2}.$
Proof.
We prove the first inequality, as the second follows analogously. Again, we can assume that the operator-valued function $f$ is positive. Observe that for any sequence $(a_{k})$,
$\sum_{k=t}^{n}B_{k}^{t+1}\,(a_{k}-a_{k-1})=$ $B_{n}^{t+1}\,a_{n}-\,\sum_{k=t}^{n-1}(B_{k+1}^{t+1}-B_{k}^{t+1})\,a_{k}-B_{t}^{t+1}a_{t-1}$
$=$ $B_{n}^{t+1}\,a_{n}-(t+1)\,\sum_{k=t}^{n-1}B_{k}^{t}\,a_{k}.$
Applying above formula to $a_{k}=S_{k}^{-t}$, we deduce
(14) $B_{n}^{t+1}\,S_{n}^{-t}=\sum_{k=t}^{n}B_{k}^{t+1}\Delta S_{k}^{-t}+(t+1)\sum_{k=t}^{n-1}B_{k}^{t}S_{k}^{-t},$
where $\Delta S_{k}^{-t}=S_{k}^{-t}-S_{k-1}^{-t}$. We claim that the desired maximal inequality is equivalent to
(15) $\Big\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}\frac{1}{(1+n)^{2}}B_{n}^{t+1}S_{n}^{-t}f\Big\|_{2}\leq C_{t}\|f\|_{2}.$
Indeed, for each positive integer $n$, one has
$\frac{1}{(1+n)^{2}}B_{n}^{t+1}\leq\frac{(1+n)^{t+1}}{(1+n)^{2}}=(1+n)^{t-1}.$
On the other hand, observe that
$\frac{(1+n)^{t+1}}{B_{n}^{t+1}}=\frac{1+n}{n}\cdot\cdot\cdot\frac{1+n}{n-t}\leq\Big(\frac{1+n}{n-t}\Big)^{t+1}=\Big(1+\frac{1+t}{n-t}\Big)^{t+1}.$
If $0<t<\frac{n}{2}$, then
$\frac{1+t}{n-t}<\frac{2t+2}{n}<\frac{2t+2}{2t}\leq 2.$
If $\frac{n}{2}\leq t\leq n-1$, then $\frac{1+t}{n-t}\leq t+1$. Therefore, there is some constant $C_{t}$ depending only on $t$ such that $\frac{(1+n)^{t+1}}{B_{n}^{t+1}}\leq C_{t}$. This proves the claim.
To prove (15), we use (14) to get
$\frac{B_{n}^{t+1}}{(1+n)^{2}}\,S_{n}^{-t}f$ $=\frac{1}{(1+n)^{2}}\sum_{k=t}^{n}B_{k}^{t+1}\Delta S_{k}^{-t}f+\frac{t+1}{(1+n)^{2}}\sum_{k=t}^{n-1}B_{k}^{t}S_{k}^{-t}f$
$=:G^{1}_{n,t}f+G^{2}_{n,t}f.$
By the convexity of the operator-valued function $x\mapsto|x|^{2}$ (see e.g. [20, (1.13)]),
$|G^{1}_{n,t}f|\leq\frac{1}{(1+n)^{2}}\big(\sum_{k=t}^{n}(B_{k}^{t+1})^{2}(k+1)^{-2t+1}\big)^{\frac{1}{2}}\Big(\sum_{k=t}^{n}(k+1)^{2t-1}|\Delta S_{k}^{-t}f|^{2}\Big)^{\frac{1}{2}}.$
Observe that for $k\geq t+1$,
$B_{k}^{t+1}=k(k-1)\cdot\cdot\cdot(k-t)\leq k^{t+1},$
which gives that
$\frac{1}{(1+n)^{2}}\big(\sum_{k=t}^{n}(B_{k}^{t+1})^{2}(k+1)^{-2t+1}\big)^{\frac{1}{2}}\leq\frac{1}{(1+n)^{2}}\big(\sum_{k=t}^{n}k^{3}\big)^{\frac{1}{2}}\leq 1.$
On the other hand, it follows from [22, Lemma 2] that $\Delta S_{k}^{-t}=S_{k}^{-t-1}$. Therefore, we find
$|G^{1}_{n,t}f|\leq\Big(\sum_{k=n}^{2n}(k+1)^{2t-1}|S_{k}^{-t-1}f|^{2}\Big)^{\frac{1}{2}}\leq R^{1}_{t,m}f.$
We now turn to the second term $G^{2}_{n,t}f$. Using the convexity of the operator-valued function $x\mapsto|x|^{2}$ again, we have
$|G^{2}_{n,t}f|$ $\leq\frac{t+1}{(1+n)^{2}}\big(\sum_{k=t}^{n-1}k^{3}\big)^{\frac{1}{2}}\big(\sum_{k=t}^{n-1}k^{-3}(B_{k}^{t})^{2}|S_{k}^{-t}f|^{2}\big)^{\frac{1}{2}}$
$\leq C_{t}\big(\sum_{k=t}^{n-1}k^{2t-3}|S_{k}^{-t}f|^{2}\big)^{\frac{1}{2}}\leq C_{t}R^{1}_{t,m}f.$
Now let $(g_{n})\subset L^{+}_{2}(\mathcal{N})$ with $\|\sum_{n}g_{n}\|_{2}\leq 1$. Then by applying the estimates of $G^{1}_{n,t}f$ and $G^{2}_{n,t}f$ obtained above, we get
$\Big|\tau\Big(\frac{1}{(1+n)^{2}}B_{n}^{t+1}S_{n}^{-t}fg_{n}\Big)\Big|$ $\leq\tau(|G^{1}_{n,t}f|g_{n})+\tau(|G^{2}_{n,t}f|g_{n})$
$\leq C_{t}\tau(R^{1}_{t,m}fg_{n}).$
Hence, by Lemma 4.5,
$\Big|\sum_{n}\tau\Big(\frac{1}{(1+n)^{2}}B_{n}^{t+1}S_{n}^{-t}fg_{n}\Big)\Big|\leq C_{t}\|R^{1}_{t,m}f\|_{2}\|\sum_{n}g_{n}\|_{2}\leq C_{t}\|f\|_{2}.$
This implies (15). Thus the proof is finished. ∎
Lemma 4.6.
Let $t\in{\mathbb{N}}$ and $\beta\in{\mathbb{R}}$. Then there exists a constant $C_{t}$ such that for any $f\in L_{2}(\mathcal{N})$,
$\big\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}M_{n}^{-t+i\beta}f\big\|_{2}\leq C_{t}e^{3\beta^{2}}\|f\|_{2},$
$\big\|{\underset{0\leq n\leq\frac{d}{m+1}}{\mathrm{sup}^{+}}}N_{n}^{-t+i\beta}f\big\|_{2}\leq C_{t}e^{3\beta^{2}}\|f\|_{2}.$
Proof.
We prove the first statement, as the other one can be done similarly. Assume that $f$ is positive and $n\leq 10t$ first. By the convolution formula stated in [22, Lemma 2], one has
$S^{-t+i\beta}_{n}f=\sum_{k=0}^{n}A_{n-k}^{i\beta}S_{k}^{-t-1}f.$
Note that $|A_{n-k}^{i\beta}|\leq c_{0}e^{2\beta^{2}}$ for all $k\leq n$ by Lemma 4.1 (ii). Then it follows from the triangle inequality and Proposition 2.3 (iii) that
$\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}M_{n}^{-t+i\beta}f\|_{2}$ $\leq\sup_{0\leq k\leq n}|A_{n-k}^{i\beta}|\sum_{k=0}^{n}\frac{(1+n)^{t-1}}{(1+k)^{t}}\|{\underset{0\leq k\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}M_{k}^{-t-1}f\|_{2}$
$\leq c_{t}e^{2\beta^{2}}\|f\|_{2},$
where the second inequality holds since $n\leq 10t$ and Lemma 4.5.
Let us now analyze the case when $n>10t$. Set $n_{2}=\lfloor\frac{n}{2}\rfloor$ and consider the decomposition
$S^{-t+i\beta}_{n}f=\sum_{k=0}^{n_{2}+t}A_{n-k}^{i\beta}S_{k}^{-t-1}f+\sum_{k=n_{2}+t+1}^{n}A_{n-k}^{i\beta}S_{k}^{-t-1}f=:F_{n,t}^{1}f+F_{n,t}^{2}f.$
One can deduce from the triangle inequality and Proposition 2.3 that
$\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}M_{n}^{-t+i\beta}f\|_{2}$ $\leq\Big\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}\frac{F_{n,t}^{1}f}{(1+n)^{-t+1}}\Big\|_{2}+\Big\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}\frac{F_{n,t}^{2}f}{(1+n)^{-t+1}}\Big\|_{2}.$
We now estimate each term separately. For the second term, we use Proposition 2.3 (iii) to obtain
$\Big\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}\frac{F_{n,t}^{2}f}{(1+n)^{-t+1}}\Big\|_{2}\leq\sup_{0\leq k\leq n}|A_{n-k}^{i\beta}|\sum_{k=n_{2}+t+1}^{n}\frac{(1+n)^{t-1}}{(1+k)^{t}}\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}M_{k}^{-t-1}f\|_{2}.$
The summation can be bounded as follows
$\sum_{k=n_{2}+t+1}^{n}\frac{(1+n)^{t-1}}{(1+k)^{t}}\leq\frac{(n-n_{2}-t)(1+n)^{t-1}}{(n_{2}+t+2)^{t}}\leq\Big(\frac{n+1}{n_{2}+2}\Big)^{t-1}\leq 2^{t}.$
Combining this with Lemma 4.1 (ii) and Lemma 4.5 yields that
$\Big\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}\frac{F_{n,t}^{2}f}{(1+n)^{-t+1}}\Big\|_{2}\leq c_{t}e^{2\beta^{2}}\|f\|_{2}.$
It remains to estimate the first term. Observe that for every nonnegative integer $\ell$ (see [18, Lemma 2.4]),
(16) $\sum_{k=0}^{m}A_{n-k}^{-\ell+i\beta}S_{k}^{-t-1+\ell}f=A_{n-m}^{-\ell+i\beta}S_{m}^{-t-1+\ell}f+\sum_{k=0}^{m-1}A_{n-k}^{-\ell-1+i\beta}S_{k}^{-t+\ell}f.$
Then by repeated applications of identity (16), we obtain
$F_{n,t}^{1}f$ $=A^{i\beta}_{n-(n_{2}+t)}S_{n_{2}+t}^{-t-1}f+\sum_{k=0}^{n_{2}+t-1}A^{-1+i\beta}_{n-k}S_{k}^{-t}f$
$=A^{i\beta}_{n-(n_{2}+t)}S_{n_{2}+t}^{-t-1}f+\sum_{\ell=1}^{t-1}A^{-\ell+i\beta}_{n-(n_{2}+t-\ell)}S_{n_{2}+t-\ell}^{-t-1+\ell}f+\sum_{k=0}^{n_{2}}A^{-t+i\beta}_{n-k}S_{k}^{-1}f.$
Therefore, by Minkowski’s inequality and Proposition 2.3 (iii), we get
$\Big\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}\frac{F_{n,t}^{1}f}{(1+n)^{-t+1}}\Big\|_{2}\leq I\times II,$
where
$I=\sum_{\ell=0}^{t-1}\|{\underset{0\leq n\leq\frac{md}{m+1}}{\mathrm{sup}^{+}}}M_{n}^{-\ell}f\|_{2}$
and
$II=(1+n)^{t-1}\max\Big\{\sum_{\ell=0}^{t-1}\frac{|A^{-\ell+i\beta}_{n-(n_{2}+t-\ell)}|}{(n_{2}+t-\ell+1)^{t-\ell}},\sum_{k=0}^{n_{2}}|A^{-t+i\beta}_{n-k}|\Big\}.$
By Lemma 4.5, $I\leq C_{t}\|f\|_{2}$. On the other hand, for each $0\leq\ell\leq t-1$, we use Lemma 4.1 to get
$\frac{(1+n)^{t-1}|A^{-\ell+i\beta}_{n-(n_{2}+t-\ell)}|}{(n_{2}+t-\ell+1)^{t-\ell}}$ $\leq\frac{(1+n)^{t-1}a_{0}e^{3\beta^{2}}(n-(n_{2}+t-\ell)+1)^{-\ell}}{(n_{2}+t-\ell+1)^{t-\ell}}$
$\leq c_{t,\ell}e^{3\beta^{2}};$
and the sum
$(1+n)^{t-1}|\sum_{k=0}^{n_{2}}A_{n-k}^{-t+i\beta}|$ $\leq(1+n)^{t-1}(n_{2}+1)\cdot\max_{0\leq k\leq n_{2}}(n-k+1)^{-t}\cdot B_{t}e^{3\beta^{2}}$
$\leq c_{t}e^{3\beta^{2}}.$
Combining these two estimates, we have $II\leq c_{t}e^{3\beta^{2}},$ which finishes the proof. ∎
Now we are in a position to prove Theorem 4.2. Our approach is based on Stein’s complex interpolation. For the sake of self-completeness, we present a detailed proof.
Proof of Theorem 4.2..
Write $\alpha=\beta+i\gamma$ with $\beta,\gamma\in{\mathbb{R}}$. Choose $\theta\in(0,1),\;q\in(1,\infty)$, $a\in{\mathbb{Z}}$ and $b>0$ such that
$\frac{1}{p}=\frac{1-\theta}{2}+\frac{\theta}{q}\quad\mbox{and}\quad\beta=(1-\theta)a+\theta\,b.$
Let $f\in L_{p}(\mathcal{N})$ and $g=(g_{n})$ be a finite sequence in $L_{p^{\prime}}(\mathcal{N})$ such that $\|f\|_{p}<1$ and $\|g\|_{L_{p^{\prime}}(\mathcal{N};\ell_{1})}<1$. Define
$h(z)=u\,|f|^{\frac{p(1-z)}{2}+\frac{pz}{q}}\,,\quad z\in{\mathbb{C}},$
where $f=u|f|$ is the polar decomposition of $f$. By construction, for any $t\in\mathbb{R}$, $\|h(it)\|_{2}=\|f\|_{p}^{\frac{p}{2}}<1$ and $\|h(1+it)\|_{q}=\|f\|_{p}^{\frac{p}{q}}<1$. By Proposition 2.3 (iv), there is a function $w=(w_{n})_{n}$ continuous on the strip $\{z\in{\mathbb{C}}:0\leq{\rm Re}(z)\leq 1\}$ and analytic in the interior such that $w(\theta)=g$ and
$\sup_{t\in{\mathbb{R}}}\,\max\Big\{\|w(i\,t)\|_{L_{2}(\mathcal{N};\ell_{1})}\ ,\ \|w(1+i\,t)\|_{L_{q^{\prime}}(\mathcal{N};\ell_{1})}\Big\}<1.$
Now define
$F(z)=\exp\big(\delta(z^{2}-\theta^{2}))\,\sum_{n}\varphi\big[M_{n}^{(1-z)a+zb+i\,\gamma}(h(z))\,w_{n}(z)\big],$
where $\delta>0$ is a constant to be specified later. Clearly, $F$ is a function analytic in the open strip $\{z\in{\mathbb{C}}:0<{\rm Re}(z)<1\}$ (note that $F$ is a finite sum). Applying Lemma 4.3 when $a\geq 0$ and Lemma 4.6 when $a\leq-1$, we have
$|F(it)|$ $\leq$ $\exp\big(\delta(-t^{2}-\theta^{2}))\,\big\|\big(M_{n}^{a+i(-ta+tb+\gamma)}(h(it))\big)_{n}\big\|_{L_{2}(\mathcal{N};\ell_{\infty})}\,\big\|w(i\,t)\big\|_{L_{2}(\mathcal{N};\ell_{1})}$
$\leq$ $C_{\alpha}\,\exp\big((-\delta+c_{m,\beta,b,\gamma})t^{2}-\delta\theta^{2})\,\|h(it)\|_{2}$
$\leq$ $C_{\alpha}\,\exp\big((-\delta+c_{m,\beta,b,\gamma})t^{2}-\delta\theta^{2})\ .$
Similarly, by Lemma 4.3,
$|F(1+it)|\leq C_{\alpha,q}\,\exp\big((-\delta+c^{\prime}_{m,\beta,b,\gamma})t^{2}+\delta(1-\theta^{2}))\ .$
Choosing $\delta>\max(c_{m,\beta,b,\gamma},\;c^{\prime}_{m,\beta,b,\gamma})$, we get
$\sup_{t\in{\mathbb{R}}}\,\max\Big\{|F(i\,t)|\ ,\ |F(1+i\,t)|\Big\}\leq C_{m,\alpha,p}$
for some constant $C_{m,\alpha,p}>0$. Then by the maximum principle, $|F(\theta)|\leq C_{m,p,\beta,b,\gamma},$ which by duality implies
$\big|\sum_{n}\varphi\big[M_{n}^{\alpha}(f)\,w_{n}\big]\big|\leq C_{m,\alpha,p}\ ,$
as $g=(g_{n})$ is an abitrary finite sequence such that $\|g\|_{L_{p^{\prime}}(\mathcal{N};\ell_{1})}<1$. This proves the first inequality. The second conclusion follows similarly using the estimates involving $N_{n}^{\alpha}(f)$ from the previous lemmas. The details are left to readers. ∎
5. Proof of Theorem 1.3
In this section, we prove Theorem 1.3.
Proof of Theorem 1.3..
Let $({\mathcal{M}},\tau)$ be a semi-finite von Neumann algebra and $\alpha:{\mathbb{Z}}_{m+1}^{d}\to\text{Aut}(\mathcal{M},\tau)$ be a $\tau$-preserving action. Then for every $s\in{\mathbb{Z}}^{d}_{m+1}$, the map $\alpha_{s}$ extends to a positive isometry on $L_{p}({\mathcal{M}})$ for all $p\geq 1$. It follows from [12, Proposition 5.3] that $(\alpha_{s}\otimes id_{\ell_{\infty}})_{s\in{\mathbb{Z}}^{d}_{m+1}}$ extends to a family of isometry on $L_{p}({\mathcal{M}};\ell_{\infty})$ for $1\leq p\leq\infty$, still denoted by $(\alpha_{s})_{s}$. Hence, for each $s\in{\mathbb{Z}}^{d}_{m+1}$,
$\big\|(M_{k}x)_{1\leq k\leq d}\big\|_{L_{p}({\mathcal{M}};\ell_{\infty})}=\big\|(\alpha_{-s}M_{k}x)_{1\leq k\leq d}\big\|_{L_{p}({\mathcal{M}};\ell_{\infty})},$
which implies that
(17) $\big\|(M_{k}x)_{1\leq k\leq d}\big\|^{p}_{L_{p}({\mathcal{M}};\ell_{\infty})}=\frac{1}{|{\mathbb{Z}}^{d}_{m+1}|}\sum_{s\in{\mathbb{Z}}^{d}_{m+1}}\big\|(\alpha_{-s}M_{k}x)_{1\leq k\leq d}\big\|_{L_{p}({\mathcal{M}};\ell_{\infty})}^{p}.$
Given $x$, define an operator-valued function $f\in L_{p}({\mathbb{Z}}^{d}_{m+1},L_{p}(\mathcal{M}))$ as
$f(s)=\alpha_{-s}x,\quad s\in{\mathbb{Z}}^{d}_{m+1}.$
Then clearly for each $s\in{\mathbb{Z}}^{d}_{m+1}$,
(18) $\alpha_{-s}M_{k}x=\frac{1}{\binom{d}{k}m^{k}}\sum_{u:|u|=k}\alpha_{u-s}x=\frac{1}{\binom{d}{k}m^{k}}\sum_{u:|u|=k}f(s-u)=T_{k}f(s).$
By the definition of $\bigl\|(T_{k}f)_{1\leq k\leq d}\bigr\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}$, there exists a factorization $T_{k}f=ay_{k}b$ with $a,b\in L_{2p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}})$ and $y_{k}\in L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}}$ such that
$\|a\|_{2p}\sup_{1\leq k\leq d}\|y_{k}\|_{\infty}\|b\|_{2p}\leq\bigl\|(T_{k}{f})_{1\leq k\leq d}\bigr\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}+\delta.$
By Hölder’s inequality, one has
$\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\big\|(T_{k}{f}(s))_{1\leq k\leq d}\big\|_{L_{p}({\mathcal{M}};\ell_{\infty})}^{p}$
$\leq$ $\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\big\|a(s)\big\|_{2p}^{p}\sup_{1\leq k\leq d}\big\|y_{k}(s)\big\|_{\infty}^{p}\big\|b(s)\big\|_{2p}^{p}$
$\leq$ $\big(\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\big\|a(s)\big\|_{2p}^{2p}\big)^{\frac{1}{2}}\sup_{1\leq k\leq d}\sup_{s\in{\mathbb{Z}}_{m+1}^{d}}\big\|y_{k}(s)\big\|_{\infty}^{p}\big(\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\big\|b(s)\big\|_{2p}^{2p}\big)^{\frac{1}{2}}$
$\leq$ $\|a\|_{2p}^{p}\sup_{1\leq k\leq d}\|y_{k}\|_{\infty}^{p}\|b\|_{2p}^{p}$
$\leq$ $\big(\|(T_{k}{f})_{1\leq k\leq d}\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}+\delta\big)^{p}.$
By the arbitrariness of $\delta$, we obtain
(19) $\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\|(T_{k}{f}(s))_{1\leq k\leq d}\|_{L_{p}({\mathcal{M}};\ell_{\infty})}^{p}\leq\|(T_{k}{f})_{1\leq k\leq d}\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}^{p}.$
Now using (17), (18) and (19), we deduce that
$\big\|(M_{k}x)_{1\leq k\leq d}\big\|^{p}_{L_{p}({\mathcal{M}};\ell_{\infty})}$ $=\frac{1}{|{\mathbb{Z}}^{d}_{m+1}|}\sum_{s\in{\mathbb{Z}}^{d}_{m+1}}\big\|(T_{k}f(s))_{1\leq k\leq d}\big\|_{L_{p}({\mathcal{M}};\ell_{\infty})}^{p}$
$\leq\frac{1}{|{\mathbb{Z}}^{d}_{m+1}|}\|(T_{k}f)_{1\leq k\leq d}\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}^{p}$
$\leq C_{m,p}\frac{1}{|{\mathbb{Z}}^{d}_{m+1}|}\|f\|_{L_{p}({\mathbb{Z}}_{m+1}^{d};L_{p}({\mathcal{M}}))}^{p}$
$=C_{m,p}\frac{1}{|{\mathbb{Z}}^{d}_{m+1}|}\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\|\alpha_{s}x\|_{L_{p}({\mathcal{M}})}^{p}.$
$=C_{m,p}\|x\|_{L_{p}({\mathcal{M}})}^{p}.$
where the second inequality above uses the strong type $(p,p)$ estimate of $(T_{k})_{1\leq k\leq d}$ by Theorem 1.1, and the last equality uses the fact that $\alpha_{s}$ is isometric on $L_{p}({\mathcal{M}})$ for each $s\in{\mathbb{Z}}_{m+1}^{d}$. This proves the first estimate.
For the second assertions, let $x\in L_{p}({\mathcal{M}})$. Note that for each $k$, $M_{k}$ is unital and completely positive. Then by Kadison-Schwarz inequality for unital $2$-positive map, we have $M_{k}(x^{*}x)\leq M_{k}(x)^{*}M_{k}(x).$ Now applying the first assertion to $x^{*}x$, we deduce from Remark 2.1 that there exists positive $b\in L_{p/2}({\mathcal{M}})$ such that
$\|b\|_{p/2}\leq C_{p,m}\|x^{*}x\|_{p/2}=C_{p,m}\|x\|_{p}^{2}\ \mbox{and}\ M_{k}(x^{*}x)\leq b,\ \forall k.$
Hence, $M_{k}(x)^{*}M_{k}(x)\leq M_{k}(x^{*}x)\leq b$, which implies that for each $k$, there is a $y_{k}\in{\mathcal{M}}$ such that $M_{k}(x)=y_{k}b^{1/2}$ and $\|y_{k}\|_{\infty}$. This gives the desired factorization of $(M_{k}(x))_{k}$ as an element in $L_{p}({\mathcal{M}};\ell^{c}_{\infty})$ and the proof is finished. ∎
Remark 5.1.
The above transference method actually shows that Theorem 1.3 for the fully noncommutative setting and Theorem 1.1 for the operator-valued setting are equivalent with identical constants.
Acknowledgement: L. Gao is supported by National Natural Science Foundation of China (Grant No. 12401163). B. Xu is supported by National Natural Science Foundation of China (Grant No. 12501172), and Fundamental Research Funds for the Central Universities (Grant No. 20720250061).